OpenStudy (anonymous):

the monthly cost (in dollars) of water use (in dollars) is a linear function of the amount of water used ( in hundreds of cubic feet, HCF). The cost for using 19 HCF of water is $33.56, and the cost for using 31 HCF is $49.76. What is the cost for using 22 HCF of water?

OpenStudy (anonymous):

recall that (Y2-Y1)/(X2-X1)=slope. Because of this we can interpolate. (X1,Y1) and (X2,Y2) are the bounds to the domain and range. This being said, (Y2-Y1)/(X2-X1)=slope=(Y3-Y1)/(X3-X1), where (X3,Y3) is the center point. Your your example, Let X=HCF and Y=$. You have 3 points, ((X1,Y1) and (X2,Y2) and (X3, ???). So you'll have 1 equation with 1 unknown.

OpenStudy (anonymous):

huh?